# First Astrometric Limits on Binary Planets and Exomoons orbiting Pictoris b

### v1

@@ 5.3 Prospects for Improved Exomoon Searches in the Future
It will also be possible to probe a larger parameter space by combining astrometric observations with other techniques, in particular radial velocities (e.g. Vanderburg et al., 2018; Lazzoni et al., 2022). Even though both radial velocities and astrometry probe the “wobble” of the planet due to the moon, the signal amplitude of an astrometric orbit scales proportionally to the moon’s semimajor axis , while the amplitude of a radial velocity orbit scales with . Therefore, radial velocities are much more sensitive to short-period moons, while astrometry is more sensitive to long-period moons. Recently, Kenworthy et al. in prep reported the first constraints on exomoons around Pic b using radial velocity observations, achieving 350 precision and demonstrating sensitivity to sub-Saturn mass moons at short orbital periods. Combining these short-period radial velocity constraints with our strong astrometric constraints on long-period moons is a promising path forward to searching for moons at all orbital separations around directly imaged planets.
@@ 6 Conclusions
These are the first astrometric limits on the presence of exomoons and binary planets in the Pictoris system, and to our knowledge, along with Kral et al. (2025), one of the first astrometric constraints on companions around any directly imaged planet. We find that existing observations are primarily sensitive to systems where the center of mass is outside of the planet itself, which is a commonly cited criterion for distinguishing binary planets from moons, (Stern & Levison, 2002). However, these observations are sufficient to achieve sensitivity to moon/planet mass ratios smaller than that of the Earth and its moon, and increasing the number of GRAVITY observations could plausibly extend sensitivity to companions that are unambiguously moons.
@@ A.1 Likelihood function
As in Section 3.2, we define to be the correlation coefficient between astrometric measurements in the RA direction and the DEC direction for observation . Again, and refer to our astrometric measurements in those directions, while and refer to our model astrometric measurements in those directions.

### v2

@@ V.3 Prospects for Improved Exomoon Searches in the Future
It will also be possible to probe a larger parameter space by combining astrometric observations with other techniques, in particular radial velocities (e.g. Vanderburg et al., 2018; Lazzoni et al., 2022; Ruffio et al., 2023). Even though both radial velocities and astrometry probe the “wobble” of the planet due to the moon, the signal amplitude of an astrometric orbit scales proportionally to the moon’s semimajor axis , while the amplitude of a radial velocity orbit scales with . Therefore, radial velocities are much more sensitive to short-period moons, while astrometry is more sensitive to long-period moons. Recently, Kenworthy et al. in prep reported the first constraints on exomoons around Pic b using radial velocity observations, achieving 350 precision and demonstrating sensitivity to sub-Saturn mass moons at short orbital periods. Overall, combining these short-period radial velocity constraints with strong astrometric constraints on long-period moons is a promising path forward to searching for moons at all orbital separations around directly imaged planets. This will be especially important as the next generation of high-resolution spectrographs on 30m class telescopes could achieve radial velocity precisions of order 10-40 m s-1, sufficient sensitivity to detect moons with mass ratios as low as at short orbital periods (Ruffio et al., 2023).
@@ VI Conclusions
These are the first astrometric limits on the presence of exomoons and binary planets in the Pictoris system, and to our knowledge, along with Kral et al. (2025), one of the first astrometric constraints on companions around any directly imaged planet. We find that existing observations are primarily sensitive to systems where the center of mass is outside of the planet itself, which is a commonly cited criterion for distinguishing binary planets from moons, (Stern and Levison, 2002). However, these observations are sufficient to achieve sensitivity to moon/planet mass ratios smaller than that of the Earth and its moon, and increasing the number of GRAVITY observations could plausibly extend sensitivity to companions that are unambiguously moons.
@@ A.1 Likelihood function
As in Section III.2, we define to be the correlation coefficient between astrometric measurements in the RA direction and the DEC direction for observation . Again, and refer to our astrometric measurements in those directions, while and refer to our model astrometric measurements in those directions.